OpenStudy (ujjwal):

During an adiabatic process, pressure of gas is found to be directly proportional to cube of its absolute temperature. The ratio of \(C_P/C_V\) is a)2 b)5/3 c)3/2 d)4/3

OpenStudy (foolaroundmath):

For an adiabatic process, \[PV^{\gamma} = const \Rightarrow P(\frac{T}{P}) ^{\gamma} = const \Rightarrow P = cT^{\gamma/(\gamma -1)}\] comparing with given data, \[\gamma/(\gamma-1) = 3 \Rightarrow \gamma = 3/2\]

OpenStudy (ujjwal):

@FoolAroundMath , Can we replace V by (T/P)? I didn't get that part..

OpenStudy (foolaroundmath):

PV = nRT. V =(T/P)*(nR) .... replace nR = const, it will just get factored in the "const"

OpenStudy (ujjwal):

Thanks @FoolAroundMath !